# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

### v1

@@ 6.2. Regularity of $\mathcal{F}_{\nicefrac{{3}}{{2}}}$
for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$, we derive that $\gamma(u_{n},x)<\nicefrac{{3}}{{2}}$ for $n$ sufficiently large for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. On the other hand, from Theorem 1.2 we obtain that $\gamma(u_{n},x)=1$ for all $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. As a consequence, for some $n$ large enough, we can find $t>0$ such that in the set
@@
Let $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and let $u_{0}$ be the $\nicefrac{{3}}{{2}}$-homogeneous blow-up of $u$ at $x_{0}$, which is unique and non-zero by Proposition 6.3. Up to multiplying $u$ with a constant, we can suppose that $u_{0}=Y$. By Lemma 6.5 and Theorem 1.2 we know that in a neighborhood of $x_{0}$ there are only points of frequency $1$ or $\nicefrac{{3}}{{2}}$. By the no-holes lemma Lemma 6.6, we know that for every $x^{\prime\prime}\in\mathbb{R}^{d-2}$ with $|x^{\prime\prime}|$ small enough, there exist $(x_{d-1},x_{d})\in\mathbb{R}^{2}$ such that $x=(x^{\prime\prime},x_{d-1},x_{d})\in\nicefrac{{3}}{{2}}$. By Proposition 6.3, we know that the point $(x_{d-1},x_{d})$ is unique, so in a neighborhood of $x_{0}$, $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ is a graph of a function $\eta:\mathbb{R}^{d-2}\to\mathbb{R}^{2}$. The $C^{1,\alpha}$-regularity of $\eta$ follows from Proposition 6.3, Proposition 6.4 and Lemma 6.1 by a standard argument. Finally, the regularity of $\mathcal{F}(u)$ around $x_{0}$ is a consequence of the regularity of $\eta$, the uniqueness of the blow-up and the epsilon-regularity for $\mathcal{F}_{1}$ ([OV24, Lemma 9.8]); the argument is standard and for the details we refer for instance to [OV24, Section 9]. ∎

### v2

@@ 6.2. Regularity of $\mathcal{F}_{\nicefrac{{3}}{{2}}}$
for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$, we derive that $\gamma(u_{n},x)<\nicefrac{{3}}{{2}}$ for $n$ sufficiently large for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. On the other hand, by the optimal frequency gap we also know that $\gamma(u_{n},x)=1$ for all $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. As a consequence, for some $n$ large enough, we can find $t>0$ such that in the set
@@
Let $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and let $u_{0}$ be the $\nicefrac{{3}}{{2}}$-homogeneous blow-up of $u$ at $x_{0}$, which is unique and non-zero by Proposition 6.3. Up to multiplying $u$ with a constant, we can suppose that $u_{0}=Y$. By Lemma 6.5 we know that in a neighborhood of $x_{0}$ there are only points of frequency $1$ or $\nicefrac{{3}}{{2}}$. By the no-holes lemma Lemma 6.6, we know that for every $x^{\prime\prime}\in\mathbb{R}^{d-2}$ with $|x^{\prime\prime}|$ small enough, there exist $(x_{d-1},x_{d})\in\mathbb{R}^{2}$ such that $x=(x^{\prime\prime},x_{d-1},x_{d})\in{\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)}$. By Proposition 6.3, we know that the point $(x_{d-1},x_{d})$ is unique, so in a neighborhood of $x_{0}$, $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ is a graph of a function $\eta:\mathbb{R}^{d-2}\to\mathbb{R}^{2}$. The $C^{1,\alpha}$-regularity of $\eta$ follows from Proposition 6.3, Proposition 6.4 and Lemma 6.1 by a standard argument. Finally, the regularity of $\mathcal{F}(u)$ around $x_{0}$ is a consequence of the regularity of $\eta$, the uniqueness of the blow-up and the epsilon-regularity for $\mathcal{F}_{1}{(u)}$ ([OV24, Lemma 9.8]); the argument is standard and for the details we refer for instance to [OV24, Section 9]. ∎